OpenStudy (anonymous):

Solve the problem. Use π= 3.14 when necessary. Find the volume of a rectangular pyramid with base 17 m2 and height 7 m. Find the result to the nearest unit. A. 38 m3 B. 119 m3 C. 60 m3 D. 40 m3

jimthompson5910 (jim_thompson5910):

V = (Area of base*Height)/3 V = (17*7)/3 V = 119/3 V = 39.666666666666 which rounds to 40 cubic meters

OpenStudy (anonymous):

thanx

jimthompson5910 (jim_thompson5910):

yw